The number of programming errors in HW assignments by computer science students is assumed to follow a Poisson distribution with mean l (so that the density function is f(x) = lambda ^ x * (e ^ - lambda) / x! ).

a) Recall that the mean and variance of a Poisson distribution are both equal to l. A professor examines the distribution of the number of programming errors for one of her assignments. There were 50 assignments turned in. Suppose that lambda = 5 in this case. What are the mean and variance of Xbar?

b) What does the central limit theorem tell us about the distribution of Xbar?

c) Assuming that n=50 is large enough for the CLT to be valid, what is the probability that the average number of errors is 5.0 or higher? What is the probability that the average number of errors is 6.0 or higher?

d) In past years the average number of errors for this assignment has been 5.0. This year the average is 6.0. The teaching assistant argues that the class is just as good, or better, as past years and the higher average is just due to chance. What do you think? Justify your answer. (Hint: find the p-value for H0: lambda = 5 vs H1: lambda > 5)
Answer #1

Answer: number of Programming Let a denotes the errors in HW asslagments is assumed to follow a Poisson distribution with meac) By using CLT, INN(5,5/6) ; where xn Poisson (5) mean u=5 n=5 XNN (5,410) P (735) = P(ante = P(230) 30.5 5-5 P (X>6) = P(12Test statistics : to 6-5 1515150 T50 Vड =5103.16 P-value = P(Z >3-16) = 0.00079. critical value 70.05 - 1.645 and at a = 0.05